OpenStudy (anonymous):

Simplify, assuming that the factors are approximations. Give answers in scientific notation with the same number of significant digits as in t he least accurate factor. (8 x 10^6) x (2 x 10^-2) / 4 x 10^2

OpenStudy (anonymous):

Please anyone help!

jhonyy9 (jhonyy9):

8*10^6 * ( 2* (1/10^2)) --------------------- = 2*10^4 * 2/10^2 = 4*10^2 4*10^2

jhonyy9 (jhonyy9):

do you understand it sure ?

OpenStudy (anonymous):

@jhonyy9 No i dont :( can you explain?

jhonyy9 (jhonyy9):

not is correct right ?

jhonyy9 (jhonyy9):

what you not can understanding ?

OpenStudy (anonymous):

@jhonyy9 I dont understand how you simplified it!

jhonyy9 (jhonyy9):

so first hope so much that you know sure that a^(-1) = 1/a yes ? so than in your exercise 10^(-2) = 1/(10^2) ok ?

jhonyy9 (jhonyy9):

so hence inside second parentheses will be 2 ------- 10^2 ok ?

jhonyy9 (jhonyy9):

so after this there are 10^6 / 10^2 = 10^(6-2) =10^4 yes ?

jhonyy9 (jhonyy9):

so again there are 2 2*10^4 * ----------- = 4*10^(4-2) = 4*10^2 10^2 so hope so much that now is understandably sure right easy good luck bye

OpenStudy (anonymous):

Yes! Thank you very much @jhonyy9

jhonyy9 (jhonyy9):

was my pleasure bye